# Characterizing and Efficiently Accelerating Multimodal Generation Model Inference

### v1

@@ Abstract
Generative artificial intelligence (AI) technology is revolutionizing the computing industry. Not only its applications have broadened to various sectors but also poses new system design and optimization opportunities. The technology is capable of understanding and responding in multiple modalities. However, the advanced capability currently comes with significant system resource demands. To sustainably scale generative AI capabilities to billions of users in the world, inference must be fast and efficient. This paper pinpoints key system design and optimization opportunities by characterizing a family of emerging multi-modal generation models on real systems. Auto-regressive token generation is a critical latency performance bottleneck, typically dominated by GPU idle time. In addition to memory-intensive attention across the generative AI models, linear operations constitute significant inference latency due to the feed forward networks in Transformer-based models. We demonstrate that state-of-the-art optimization levers, spanning from applications to system software and hardware, set a 3.88$\times$×better baseline.
I Introduction
Generative AI technologies are driving an unprecedented growth for the computing industry, introducing a new paradigm shift for AI. This technology redefines the interaction between humans and AI by enabling the creation of highly realistic images [38], videos [39, 14], texts, and speech [6], as well as intricate textual patterns or even new materials. Large language models (LLMs), such as ChatGPT [28], Llama [44, 45], or Gemini [43], demonstrate remarkable capabilities. LLMs not only enhance user experience by providing contextually relevant interactions but also play a critical role in automating complex tasks. It has already germinated a wide variety of applications, leading to higher productivity.
Beyond LLMs, multi-lingual speech translation and transcription models, such as Seamless [6], Whisper [33], or Translatotron [22], are pivotal in breaking down the language barriers and enhancing communication on a global scale. The speech models provide accurate and real-time translation/transcription across different languages by processing speech and text modalities together, such as Speech to Speech and Text (S-ST), Text to Speech and Text (T-ST), and Automatic Speech Recognition (ASR).
In addition to text and speech modalities, state-of-the-art AI technologies can take inputs of multiple modalities to serve multi-modal use cases. Taking Chameleon [42] as an example, this multi-modal foundation model can take images and text as input and generate outputs in either modality. Such models are the foundation of image editing or visual question-answer (VQA) use cases. Also, the multi-modal models are capable of image generation based on text prompts or even ChatBot style conversations.
@@ I Introduction
Beyond learning from texts, language, speeches, images or videos, generative AI technologies are also adopted in deep learning recommendation systems as well. Leveraging the ability of Attention-based Transformers for automatically extracting and learning features from datasets, recent deep learning recommendation models, such as HSTU [50], TIGER [34], introduce a new feature generation paradigm by adopting sequential generative models. Such new model architecture uses generative models to accurately predict items of interest. Generative recommendation models overcome the model quality saturation problem faced by existing deep learning recommendation models (DLRMs) [25], exceeding prediction quality over prior recommender system technologies.
While a disproportional investment is currently focused on LLMs, generative AI technologies that are capable of processing multi-modal inputs and outputs are on the horizon. Depending on distributions of input prompt lengths and use cases (Section III-A) and the characteristics of model architectures (Section III-B), the system design space for efficiency presents unique optimization opportunities. For example, a recent work shows that training a state-of-the-art text-to-image model can use 14x more GPUs per model parameter than that of an industry-scale LLM [15]. To efficiently accelerate multi-modal generation model inference, this paper provides an in-depth system performance characterization for important industry-scale generative AI tasks: language (Code Llama [35]), speech translation (Seamless [6]), text and image generation (Chameleon [42]), and generative deep learning recommender systems (gDLRM [50]).
@@
Figure 2: Model Architectures of Llama [35], Seamless [6], Chameleon [42], HSTU [50].
@@
To sustainably scale generative AI technologies for a large, diverse variety of applications [47], we must understand and enable AI deployment in a resource-efficient manner [49]. Figure 1 illustrates the system requirements for four multi-modal generation models at a single batch size. The chart highlights the latency requirement, overall memory capacity, communication requirement, and GPU utilization, for different tasks across these models. It is evident that, depending on input modalities and model architectures of a specific task, system resource utilization characteristics are distinct. For example, Chameleon can perform image-to-text (I-T), text-to-image (T-I), and image-text-to-text (IT-T) without requiring fine-tuning for each task. However, the T-I task demands significantly higher resource requirements across all four axes.
@@
To scale advanced generative AI capabilities to billions of user in the world, inference needs to complete in the order of millisecond and efficiently. The in-depth real-system performance characterization results in Section III-B guide the focus of inference performance and efficiency optimization. We take a step further to enable state-of-the-art inference performance optimization techniques — torch.compile and CUDA Graph for memory efficiency optimization [1], Scaled Dot Product Attention (SDPA) / Flash Attention to speed up Transformer’s key performance bottleneck [9], quantization to further improve compute density and memory bandwidth utilization. When enabling state-of-the-art optimization levers properly, the inference performance over the important generative AI tasks can be improved by 3.8x, setting a new, more rigorous baseline. Beyond efficiently accelerating inference performance horizontally across the key generative AI tasks, in Section IV-C, we present ways to further improve inference performance efficiency with application-specific, algorithmic optimization. We enable LayerSkip [12], a self-speculative decoding approach to our workloads to speedup generation and show the inference performance is improved by 1.58$\times$.
The key contributions of this paper are as follows:
@@
  • System Performance Characterization for Emerging Multi-Modal Generative AI Tasks This paper delivers an in-depth examination of system performance across four pivotal generative AI models: LLM (Code Llama), Speech Translation (Seamless), Generative Text and Image Models (Chameleon), and Generative Deep Learning Recommendation Models (gDLRM). Our analysis covers critical aspects, such as computational demands, memory bandwidth requirements, and variations in input distributions — key to inference performance efficiency optimization.
@@
  • Optimized Baseline for Generative AI Inference Acceleration We demonstrate the importance of enabling state-of-the-art optimization methods — torch.compile, CUDA Graph, SDPA/Flash Attention, and quantization — that accelerate the inference performance across the generative AI tasks by upto 28$\times$. Algorithmic optimization — LayerSkip — improves inference performance as well by 1.58$\times$. Altogether, cross-stack solutions, spanning algorithm and systems, improve inference performance by an average of 3.88$\times$.
@@
  • Design Implications and New Directions for Future Systems We distill the implications of our findings for future research and development — 1) New solutions must improve upon stronger baseline 2) With proper understandings of the distinct characteristics and end-to-end inference pipeline of a given model, we can achieve 3.88$\times$ speedup with state-of-the-art optimizations leverages 3) Enhancing the baseline with software optimization methods unlocks new possibilities for both current and future hardware architectures.
II Background and Motivation
II-A Understanding the Lay of the Land for Multi-modal Generative AI Tasks
@@
We provide an overview for key generative AI technologies. Figure 2 illustrates the model architectures for four generative AI models — LLM (Code Llama), Speech Translation (Seamless), Generative Text and Image Models (Chameleon) and Generative Deep Learning Recommendation Models (gDLRM). Table III summarizes input/output modalities and sequence length distribution for different
@@
Text-based LLM Llama [35] T-T Code Completion, Infilling, Instruction Text Text
Image&Text Generation Chameleon [42] I-T Image Captioning Image Text
@@
IT-T Visual Question Answering Image & Text Text
@@
Speech&Text Translation Seamless [6] $\bigtriangleup$ Only text decoder S-S Speech-to-Speech Translation Speech Speech
S-T Speech-to-Text Translation Speech Text
T-T Text-to-Text Translation Text Text
T-S Text-to-Speech Translation Text Speech
@@
Generative DLRM HSTU [50] H-A Ranking and Retrieval User History Engagement Type (ranking) Recommended Item (retrieval)
@@
TABLE I: The input and output modality of each task performed by four multimodal generative models, LLM (Llama), speech&text translation (Seamless), text&image generation (Chameleon) and generative DLRM (HSTU).
II-A1 Llama for Language Generation
@@
Code Llama is a family of large language models based on Llama-2 for coding tasks. Code Llama takes text-based code as input and generates code as output. The input sequence length distribution for Code Llama can vary depending on the specific model and task it is being used for. However, in general, Code Llama models are trained on a wide range of input sequence lengths to be able to handle varying sizes of code snippets. For example, Code Llama support sequence lengths up to 100,000 tokens which is enough to capture a reasonably sized code snippet while keeping the computation feasible.
@@
The model architecture of Code Llama is a standard transformer architecture [46] as shown in Figure 2(a). In this paper, we take Code Llama as our model representing language generative AI model. And we refer Code Llama as Llama from now on for convenience. The model consists of embedding layer followed by consecutive Transformer decoder blocks that includes attention and feed forward layer. Specifically, Llama 34B model has 48 layers of Transformer decoder blocks.
@@
Llama is an autoregressive generation model where inference pipeline is broken down into two phases: prefill (Prefill) and incremental decoding (Decoding) . Prefill takes the full input prompt whereas Decoding generates output tokens one by one based on previously generated tokens. In Decoding, KV cache optimization is key to relieve computational intensity. The input sequence length for Decoding is always 1 with KV cache optimization while the input sequence length for Prefill is determined by the length of the full input prompt.
II-A2 Seamless for Speech Translation
@@
Seamless[6] is a family of speech translation models that enable more natural and authentic communication across languages. SeamlessM4T is the foundation model for multilingual multimodal machine translation supporting around 100 languages. SeamlessM4T achieves state-of-the-art semantic accuracy, supports a wide range of languages, and provides multitasking capabilities from and into text or speech.
@@
SeamlessM4T, which we refer to as Seamless in this paper, consists of multiple pretrained blocks that are finetuned as a unified model. The four main building blocks are shown in Figure 2(b), including
@@
  • Conformer Speech Encoder (Blue) Conformer speech encoder is a speech representation learning model that leverages unlabeled speech audio data.
@@
  • Text-to-Text Translator (T2TT) (Pink) T2TT is a text-to-text translation model pre-trained on NLLB data in nearly 100 languages. This is the ”only” autoregressive module among all modules in Seamless.
@@
  • Vocoder (Orange) Vocoder is a HiFi-GAN unit-vocoder that converts generated units to waveform output where an unit is a representation of speech that combines different aspects such as phonemes and syllables, which can be used to generate sounds that are audible to humans.
@@
II-A3 Chameleon for Text and Image Generation
Chameleon is a foundational model for the family of early-fusion token-based mixed-modal models capable of understanding and generating images and text in any arbitrary sequence. The model is capable of performing various tasks including visual question answering, image captioning, text generation, image generation, and long-form mixed modal generation. Chameleon demonstrates broad and general capabilities, including state-of-the-art performance in image captioning tasks all in a single model. The model architecture of Chameleon largely follows Llama-2 [45] as shown in Figure 2(c), thus Chameleon is also an auto-regressive generation model. For normalization, Chameleon continue to use RMSNorm [51]; and Chameleon uses the SwiGLU [37] activation function and rotary positional embeddings (RoPE) [41]. Chameleon 34B model has 48 layers of Transformer decoder blocks.
Chameleon represents all modalities — images, text, and code, as discrete tokens and uses a uniform transformer-based architecture that is trained from scratch in an end-to-end fashion on around 10T tokens of interleaved mixed-modal data. Chameleon can take any combination of image and text and utilizes image tokenizer [13] and text tokenizer [36] respectively to generate tokens to be fed to the model. For text generation, generated tokens are decoded by text tokenizer to generate readable texts. For image generation, Chameleon generates 1024 image tokens and then detokenize them using an image detokenizer to generate images in a format that can be interpreted by human such as jpg. And Chameleon uses a contrastive decoding method specifically for T-I task, which aims to maximize the differences between a weak and a strong model. Logits from conditioned outputs are treated as the strong model, while unconditional logits are considered as the weak model. As a result, Chameleon decodes twice at each time step for T-I task.
II-A4 gDLRM for Generative Recommendations
Generative recommenders approach information retrieval and recommendation problems by modeling the underlying joint distribution of user-item interactions, and adopting homogeneous, large-scale sequential backbones to replace the traditional heterogeneous modules in DLRMs. gDLRMs enables the main tasks in recommendations, namely retrieval (predict the next item to recommend) and ranking (predict the engagement type given the retrieved item), to be formulated as a next-token prediction problem. We refer to both type of outputs for retrieval and ranking tasks as ”Action”. Compared to prior DLRMs [25, 18, 48, 21], gDLRMs have demonstrated superior accuracy performance [50] and further enable a unified feature space to be used across different domains.
One key sequential architecture used in the generative recommender system — HSTU — can be viewed as a variant of self-attention or Transformers specialized for sequence-to-sequence (sequential transduction) tasks. HSTU is composed of a stack of identical layers connected by residual connections [20] as shown in Figure 2(d). Each layer consists of three main sub-layers: Point-wise Projection, Spatial Aggregation, and Pointwise Transformation. Spatial Aggregation replaces sequence-level normalized Softmax with pointwise normalized attention and relative attention bias, and Point-wise Projection together with Pointwise Transformation together performs efficient token-level transformation augmented by element-wise gating. This reduces the number of matrix multiplication operations from standard Transformers. In general, training throughput performance can be significantly improvement through feature deduplication optimization [52]. Note that HSTU is the only model that is non-autoregressive among the generation tasks studied in this paper.
Model Dataset Modality Decoding Step Count Avg. Inf Time (ms) (5 Sample)
@@ II-A3 Chameleon for Text and Image Generation
Llama HumanEval Text 44 430 153.5 Text 55 10000 691.84 538.33 4493.66
MBPP Text 29 1748 59.41 Text 38 10000 1075.66 1016.25 5566.732
Seamless Fleurs Eng-Spa Speech 179.00 1464.00 492.83 Speech 129.00 1029.00 385.07 34.68 1577.45
Text 15 98 36.03 1320.97
Text 12 80 30.48 Speech 145 1030 393.18 33.91 1431.97
Text 14 95 35.35 1186.51
Chame- leon MSCOCO Image 1030 1030 1030 Text 30 30 30 30 2912.73
Vizwiz Img&Txt 1033 1095 1040 Text 10 10 10 10 1253.1
MSCOCO Text 10 22 13.9 Image O(1025) O(1025) O(1025) 1024 159702.4
HSTU Synthetic User History 4507.0 5121.0 4813.9 Action 4507.0 5121.0 4813.9 N/A 49.89
TABLE II: Sequence Length Distribution of Four Generative AI Models.
III System Performance Characterization on Multi-Modal Model Inference
We present real-system performance characterization results for the key generative AI tasks in this section. The deeper understanding of application-level characteristics and performance bottlenecks on real systems help guide our performance and efficiency optimization focus systematically. The data-driven analysis also underpins key system design and optimization opportunities, as what we later show in Section III-B.

### v2

@@ Abstract
Generative artificial intelligence (AI) technology is revolutionizing the computing industry, posing new system design and optimization opportunities. In particular, AI’s ability to understand and respond in multiple modalities comes with significant system resource demands. To sustainably scale generative AI capabilities to billions of users in the world, inference must be fast and efficient. This paper pinpoints key system design and optimization opportunities by characterizing a family of emerging multi-modal generation models on real systems. Auto-regressive token generation is a critical latency performance bottleneck, typically dominated by GPU idle time. In addition to memory-intensive attention across the generative AI models, linear operations constitute significant inference latency due to the feed forward networks in Transformer-based models. We demonstrate that state-of-the-art optimization levers, spanning from applications to system software and hardware, set a 3.88$\times$ better baseline.
[…]
Yejin Lee , Carole-Jean Wu
1 Introduction
Generative AI technologies are driving an unprecedented growth for the computing industry, introducing a new paradigm shift for AI. This technology redefines the interaction between humans and AI by enabling the creation of highly realistic images Sheynin et al. (2023), videos Singer et al. (2022); Girdhar et al. (2023), texts, and speech Communication et al. (2023), as well as intricate textual patterns or even new materials. Large language models (LLMs), such as ChatGPT OpenAI (2024), Llama Touvron et al. (2023a, b), or Gemini Team (2024b), demonstrate remarkable capabilities. LLMs not only enhance user experience by providing contextually relevant interactions but also play a critical role in automating complex tasks. It has already germinated a wide variety of applications, leading to higher productivity.
Beyond LLMs, multi-lingual speech translation and transcription models, such as Seamless Communication et al. (2023), Whisper Radford et al. (2022), or Translatotron Jia et al. (2019), are pivotal in breaking down the language barriers and enhancing communication on a global scale. The speech models provide accurate and real-time translation/transcription across different languages by processing speech and text modalities together, such as Speech to Speech and Text (S-ST), Text to Speech and Text (T-ST), and Automatic Speech Recognition (ASR).
In addition to text and speech modalities, state-of-the-art AI technologies can take inputs of multiple modalities to serve multi-modal use cases. Taking Chameleon Team (2024a) as an example, this multi-modal foundation model can take images and text as input and generate outputs in either modality. Such models are the foundation of image editing or visual question-answer (VQA) use cases. Also, the multi-modal models are capable of image generation based on text prompts or even ChatBot style conversations.
@@ 1 Introduction
Beyond learning from texts, language, speeches, images or videos, generative AI technologies are also adopted in deep learning recommendation systems as well. Leveraging the ability of Attention-based Transformers for automatically extracting and learning features from datasets, recent deep learning recommendation models, such as HSTU Zhai et al. (2025), TIGER Rajput et al. (2023), introduce a new feature generation paradigm by adopting sequential generative models. Such new model architecture uses generative models to accurately predict items of interest. Generative recommendation models overcome the model quality saturation problem faced by existing deep learning recommendation models (DLRMs) Naumov et al. (2019), exceeding prediction quality over prior recommender system technologies.
While a disproportional investment is currently focused on LLMs, generative AI technologies that are capable of processing multi-modal inputs and outputs are on the horizon. Depending on distributions of input prompt lengths and use cases (Section 3.1) and the characteristics of model architectures (Section 3.2), the system design space for efficiency presents unique optimization opportunities. For example, a recent work shows that training a state-of-the-art text-to-image model can use 14x more GPUs per model parameter than that of an industry-scale LLM Golden et al. (2024a). To efficiently accelerate multi-modal generation model inference, this paper provides an in-depth system performance characterization for important industry-scale generative AI tasks: language (Code Llama Rozière et al. (2024)), speech translation (Seamless Communication et al. (2023)), text and image generation (Chameleon Team (2024a)), and generative deep learning recommender systems (gDLRM Zhai et al. (2025)).
These models serve important roles in Meta’s workloads, with Llama functioning as our foundational large language model powering Meta AI’s core capabilities. Seamless delivers high-quality translation services across Meta platforms, enabling multilingual access to Instagram and Facebook content. Chameleon serves as a foundation model for multimodal generation, handling various combinations of text and image inputs/outputs, while HSTU helps drive recommendation systems, processing billions of recommendations daily with strict latency constraints.
@@
Figure 2: Model Architectures of Llama Rozière et al. (2024), Seamless Communication et al. (2023), Chameleon Team (2024a), HSTU Zhai et al. (2025).
@@
To sustainably scale generative AI technologies for a large, diverse variety of applications Wu et al. (2024), we must understand and enable AI deployment in a resource-efficient manner Wu et al. (2022). Figure 1 illustrates the system requirements for four multi-modal generation models at a single batch size. The chart highlights the latency requirement, overall memory capacity, communication requirement, and GPU utilization, for different tasks across these models. It is evident that, depending on input modalities and model architectures of a specific task, system resource utilization characteristics are distinct. For example, Chameleon can perform image-to-text (I-T), text-to-image (T-I), and image-text-to-text (IT-T) without requiring fine-tuning for each task. However, the T-I task demands significantly higher resource requirements across all four axes.
@@
To scale advanced generative AI capabilities to billions of users in the world, inference needs to complete on the order of milliseconds and efficiently. The in-depth real-system performance characterization results in Section 3.2 guide the focus of inference performance and efficiency optimization. We take a step further to enable state-of-the-art inference performance optimization techniques — torch.compile and CUDA Graph for memory efficiency optimization Ansel et al. (2024), Scaled Dot Product Attention (SDPA) / Flash Attention to speed up Transformer’s key performance bottleneck Dao et al. (2022), quantization to further improve compute density and memory bandwidth utilization. When enabling state-of-the-art optimization levers properly, the inference performance over the important generative AI tasks can be improved by 3.8x, setting a new, more rigorous baseline. Beyond efficiently accelerating inference performance horizontally across the key generative AI tasks, in Section 4.3, we present ways to further improve inference performance efficiency with application-specific, algorithmic optimization. We enable LayerSkip Elhoushi et al. (2024), a self-speculative decoding approach to our workloads to speedup generation and show the inference performance is improved by 1.58$\times$. The key contributions of this paper are as follows:
@@
  • System Performance Characterization for Emerging Multi-Modal Generative AI Tasks This paper delivers an in-depth examination of system performance across four pivotal generative AI models: LLM (Code Llama), Speech Translation (Seamless), Generative Text and Image Models (Chameleon), and Generative Deep Learning Recommendation Models (gDLRM). Our analysis covers critical aspects, such as computational and memory bandwidth requirements, variations in input distributions and roofline analysis — key to inference performance efficiency optimization.
@@
  • Optimized Baseline for Generative AI Inference Acceleration We demonstrate the importance of enabling state-of-the-art optimization methods — torch.compile, CUDA Graph, SDPA/Flash Attention, and quantization — that accelerate the inference performance across the generative AI tasks by upto 28$\times$. Algorithmic optimization — LayerSkip — improves inference performance as well by 1.58$\times$. Altogether, cross-stack solutions, spanning algorithm and systems, improve inference performance by an average of 3.88$\times$. We also highlight the performance impact of using a newer generation of GPUs by comparing the performance analysis across different GPU generation.
@@
  • Design Implications and New Directions for Future Systems We distill the implications of our findings for future research and development — 1) New solutions must improve upon stronger baseline 2) With proper understandings of the distinct characteristics and end-to-end inference pipeline of a given model, we can achieve 3.88$\times$ speedup with state-of-the-art optimizations leverages 3) Enhancing the baseline with software optimization methods unlocks new possibilities for current and future hardware architectures.
2 Background and Motivation
2.1 Understanding the Lay of the Land for Multi-modal Generative AI Tasks
@@
We provide an overview for key generative AI technologies. Figure 2 illustrates the model architectures for four generative AI models — LLM (Code Llama), Speech Translation (Seamless), Generative Text and Image Models (Chameleon) and Generative Deep Learning Recommendation Models (gDLRM). Table 1 summarizes input/output modalities and sequence length distribution for different workloads.
@@
Text-based LLM Llama T-T Code Completion, Infilling, Instruction Text Text
Image&Text Generation Chameleon I-T Image Captioning Image Text
@@
IT-T Visual Question Ans. Image & Text Text
@@
Speech&Text Translation Seamless $\bigtriangleup$ Only text decoder S-S Speech-to-Speech Trans. Speech Speech
S-T Speech-to-Text Trans. Speech Text
T-T Text-to-Text Trans. Text Text
T-S Text-to-Speech Trans. Text Speech
@@
Generative DLRM HSTU H-A Ranking and Retrieval User History Engagement Type (ranking) Recommend Item (retrieval)
@@
Table 1: The input and output modality of each task performed by four multimodal generative models, LLM (Llama), speech&text translation (Seamless), text&image generation (Chameleon) and generative DLRM (HSTU).
2.1.1 Llama for Language Generation
@@
Code Llama is a large language model for coding tasks. Code Llama models are trained on a wide range of input sequence lengths to be able to handle varying sizes of code snippets. For example, Code Llama support sequence lengths up to 100,000 tokens which is enough to capture a reasonably sized code snippet while keeping the computation feasible.
@@
Code Llama has a standard transformer architecture Vaswani et al. (2023) as shown in Figure 2(a). In this paper, we take Code Llama as our model representing language generative AI model. And we refer Code Llama as Llama from now on for convenience. The model consists of embedding layer followed by consecutive Transformer decoder blocks that includes attention and feed forward layer. Specifically, Llama 34B model has 48 layers of Transformer decoder blocks.
@@
Llama is an autoregressive generation model where inference pipeline is broken down into two phases: prefill (Prefill) and incremental decoding (Decoding). Prefill processes the entire input prompt of length $N$ at once, computing attention across all input tokens ($O(N^{2})$ complexity), whereas Decoding generates output tokens one by one based on previously generated tokens. These phases present different computational characteristics: Prefill has high computational intensity due to processing full input sequence length $N$ while Decoding is computationally lighter by using KV cache optimization that stores and reuses key-value pairs, though this makes Decoding memory-bound due to frequent cache access.
2.1.2 Chameleon for Text and Image Generation
Chameleon is a foundational model for the family of early-fusion token-based mixed-modal models capable of understanding and generating images and text. It is capable of performing broad range of tasks including visual question answering, image captioning, text and image generation, and long-form mixed modal generation in a single model. The model architecture of Chameleon largely follows Llama-2 Touvron et al. (2023b) as shown in Figure 2(b), thus Chameleon is also an auto-regressive generation model. For normalization, Chameleon continue to use RMSNorm Zhang and Sennrich (2019); and Chameleon uses the SwiGLU Shazeer (2020) activation function and rotary positional embeddings (RoPE) Su et al. (2023).
Chameleon represents images, text, and code modalities as discrete tokens and uses a uniform transformer-based architecture that is trained from scratch in an end-to-end fashion on around 10T tokens of interleaved mixed-modal data. Chameleon can take any combination of image and text and utilizes image tokenizer Gafni et al. (2022) and text tokenizer Sennrich et al. (2016) respectively to generate tokens to be fed to the model. For text generation, generated tokens are decoded by text tokenizer to generate readable texts. For image generation, Chameleon generates 1024 image tokens and then detokenize them using an image detokenizer to generate images in a format that can be interpreted by human such as jpg. And Chameleon uses a contrastive decoding method specifically for T-I task, which aims to maximize the differences between a weak and a strong model. Logits from conditioned outputs are treated as the strong model, while unconditional logits are considered as the weak model. As a result, Chameleon decodes twice at each time step for T-I task.
2.1.3 Seamless for Speech Translation
@@
Seamless Communication et al. (2023) is a family of speech translation models that enable more natural and authentic communication across languages. SeamlessM4T is the foundation model for multilingual multimodal machine translation supporting around 100 languages. SeamlessM4T achieves state-of-the-art semantic accuracy, supports a wide range of languages, and provides multitasking capabilities from and into text or speech.
@@
SeamlessM4T, which we refer to as Seamless in this paper, consists of multiple pretrained blocks that are finetuned as a unified model. The four main building blocks are shown in Figure 2(c),
@@
  • Conformer Speech Encoder (Blue) A speech representation learning model that leverages unlabeled speech audio data.
@@
  • Text-to-Text Translator (T2TT) (Pink) A text-to-text translation model pre-trained on NLLB data in nearly 100 languages. It is "only" autoregressive module among all modules in Seamless.
@@
  • Vocoder (Orange) A HiFi-GAN unit-vocoder that converts generated units to waveform output where an unit represents speech combining different aspects such as phonemes and syllables
@@
Model Dataset Modality Decode Step Count Avg. Time (ms)
@@
Llama HumanEval Text 44 430 154 Text 55 10000 692 538 4494
MBPP Text 29 1748 59 Text 38 10000 1076 1016 5567
Seam- less Fleurs Eng-Spa Speech 179 1464 493 Speech 129 1029 385 35 1578
Text 15 98 36 1321
Text 12 80 31 Speech 145 1030 393 34 1432
Text 14 95 35 1187
Chame- leon MSCOCO Image 1030 1030 1030 Text 30 30 30 30 2913
Vizwiz Img&Txt 1033 1095 1040 Text 10 10 10 10 1253
MSCOCO Text 10 22 13.9 Image O(1025) O(1025) O(1025) 1024 159702
HSTU Synthetic User History 4507 5121 4814 Action 4507.0 5121.0 4813.9 N/A 50
Table 2: Sequence Length Distribution of Four Generative AI Models. We use 5 sample for each workload.
2.1.4 gDLRM for Generative Recommendations
Generative recommenders approach information retrieval and recommendation problems by modeling the underlying joint distribution of user-item interactions, and adopting homogeneous, large-scale sequential backbones to replace the traditional heterogeneous modules in DLRMs. gDLRMs enables the main tasks in recommendations, namely retrieval (predict the next item to recommend) and ranking (predict the engagement type given the retrieved item), to be formulated as a next-token prediction problem. We refer to both type of outputs for retrieval and ranking tasks as "Action". Compared to prior DLRMs Naumov et al. (2019); Gupta et al. (2020); Wu et al. (2020); Hsia et al. (2023), gDLRMs have demonstrated superior accuracy performance Zhai et al. (2025) and further enable a unified feature space to be used across different domains.
One key sequential architecture used in the generative recommender system — HSTU — can be viewed as a variant of self-attention or Transformers specialized for sequence-to-sequence (sequential transduction) tasks. HSTU is composed of a stack of identical layers connected by residual connections He et al. (2016) as shown in Figure 2(d). Each layer consists of three main sub-layers: Point-wise Projection, Spatial Aggregation, and Pointwise Transformation. Spatial Aggregation replaces sequence-level normalized Softmax with pointwise normalized attention and relative attention bias, and Point-wise Projection together with Pointwise Transformation together performs efficient token-level transformation augmented by element-wise gating. This reduces the number of matrix multiplication operations from standard Transformers. In general, training throughput performance can be significantly improvement through feature deduplication optimization Zhao et al. (2023). Note that HSTU is the only model that is non-autoregressive among the generation tasks studied in this paper.
3 System Performance Characterization on Multi-Modal Model Inference
We present real-system performance characterization results for the key generative AI tasks in this section. The deeper understanding of application-level characteristics and performance bottlenecks on real systems help guide our performance and efficiency optimization focus systematically. The data-driven analysis also underpins key system design and optimization opportunities, as what we later show in Section 3.2.